OpenStudy (agar.io):

55/100 in simplest form @TheSmartOne I got 11/20

TheSmartOne (thesmartone):

@Agar.io @~ You are correct! :) \(\sf\Large\frac{55}{100} = \frac{11\times 5}{20\times 5} = \boxed{\frac{11}{20}}\)